OpenStudy (anonymous):

In triangle ABC, a = 50 cm, b = 24 cm, and c = 30 cm. Find the largest angle, rounded to the nearest degree.

OpenStudy (anonymous):

cosine rule. \(a^2=b^2 +c^2 -2bc cos sin A\) find A.

OpenStudy (anonymous):

u mean the law of cosine? I didnt learn that yet.

OpenStudy (anonymous):

in any case, using tangents and sines will have you end up with the cosine rule as that's how it's derived.

OpenStudy (anonymous):

So how do I approach this problem I really dont know. Like my example says, and this is a different problem btw. In right triangle ABC, a=2.73 and b = 3.41 Find the remaining side and angles. and they go on to do Tan A= a/b which equals 2.73/3.41 which equals .8006 Now they go on to say to find A, we use a calculator to look for the angle whose tangent is .8006: A=38.7 What did they do in the calculator to get 38.7?